Find all The points on the graph of (x^2+y^2+y)^2=x^2+y^2 that have horizontal tangent line


1
Expert's answer
2021-04-28T04:22:27-0400

Given, the function (x2+y2+y)2=x2+y2. ...................................(1)


First, differentiate with respect to x.

2(x2+y2+y)(2x+2yy'+y')=2x+2yy'


Second, putting y'=0.

4x(x2+y2+y)=2x

2(x2+y2+y)=1

x2+y2+y=1/2 .............................................................(2)

This is the horizontal tangent line.


Third, substitute the above horizontal tangent line in the given function.

i.e., putting equation (2) in equation (1),

(x2+y2+y)2=x2+y2

(1/2)2= -y+1/2 { from (2), x2+y2= -y+1/2 }

1/4= -y+1/2

y=1/4


Forth, finding the x-coordinate.By putting the value of y in the horizontal tangent line.

x="\pm\frac{\sqrt{3}}{4}"


Thus, the point are "(\frac{-\sqrt3}{4},\frac{1}{4})" and "(\frac{\sqrt3}{4},\frac{1}{4})" on the given graph that have horizontal tangent line.